OpenStudy (anonymous):

Use the stem-and-leaf plot of Monthly Sales Goals to answer the question that follows.

OpenStudy (anonymous):

Monthly Sales Goals (in thousands) Stem Leaf 2 4 5 8 9 3 1 3 4 6 4 5 7 5 0 0

OpenStudy (anonymous):

What is the median monthly sales goal?

OpenStudy (shamil98):

Order all numbers from least to greatest first. 0,0,1,3,4,4,5,5,6,7,8,9 the median is the value in the middle

OpenStudy (anonymous):

are you familiar with the word, 'median' it means middle number. so count in from the first number and count in from the last number on the stemLeaf Plot and see what number you end up on

OpenStudy (shamil98):

in this case you have 4 and 5 in the middle. so you add them both and divide by 2. 4 + 5 = 9 9/2 = 4.5 4.5 is your median number.

OpenStudy (anonymous):

that isnt a choice tho >_< a. $50,000 b. $33,500 c. $34,500 d. $36,000

OpenStudy (shamil98):

oh, it's in thousands.

OpenStudy (shamil98):

well, 4000 + 5000 = 9000 9000/2 = 4500 30000 + 4500 = 34500

OpenStudy (shamil98):

that's my guess atleast...

OpenStudy (texaschic101):

33000 and 34000 are your middle numbers. The median would be (33000 + 34000)/2 = 67000/2 = 33500

OpenStudy (anonymous):

okay thank you! ^o^